OpenStudy (readforlife):

Please help quick! 3. Identify the linear model you chose (A or B), justify your selection and write an equation in slope-intercept form for the linear model. Explain and identify the variables that are represented by the x-values and the y-values. 4. What are the slope and the y-intercept? Use this information to describe, with specific numbers, to new students what they might expect from attending Ms. Perez’s sessions.

OpenStudy (readforlife):

Here is the graph for it

OpenStudy (readforlife):

I already answered number 3, but I posted it cause it is necessary for question 4. Anyways, for 3. I choose Line A. I choose this line because it is a straighter line than Line B. y=mx+b (1,1) (7,8)= 8-1/7-1 = 7/6 y = 7/6x + 7 If it isn't right, please help me make it right. And I say HELP not answer. And question number 4 I am kinda confused with.

OpenStudy (readforlife):

@mathmale

OpenStudy (readforlife):

@math_man21 @mathstudyhard @mathstudent55

OpenStudy (math_man21):

ok

OpenStudy (readforlife):

Ok... So can you help....?

OpenStudy (readforlife):

@math_man21

OpenStudy (math_man21):

yes

OpenStudy (welshfella):

I think you are right for lin A because it also passes close to the origin . Line B gives one point before you do any sessions!

OpenStudy (readforlife):

Alright. Thx. And do you know what question 4 is asking? To me it looks like it wants me to solve the equation I created, but then the rest of the question I just don't get...

OpenStudy (welshfella):

Well the slope = rise / run and the y intercept is the value of y where the line crosses the yaxis

OpenStudy (welshfella):

the slope = 7/6 as you said but the y-intercept is not +7

OpenStudy (readforlife):

Oh. Ok... So... Um.... Hold on. Ima see If I can figure it out myself real quick then you can check it?

OpenStudy (welshfella):

ok

OpenStudy (readforlife):

Ok, so I put 7 which is the x-coordinate.... So would it be the y coordinate instead? y=7/6x +6??

OpenStudy (readforlife):

Wait, hold on... That's not right I don't think.

OpenStudy (welshfella):

no if we look at where the blue line crosses the y axis its about -0.5

OpenStudy (welshfella):



OpenStudy (readforlife):

Right, so either 0 or -1 because it passes through the y-intercept.

OpenStudy (readforlife):

So ima say -1?

OpenStudy (welshfella):

well you could write it as -0,5 y = (7/6)x - 0.5

OpenStudy (welshfella):

thers no reason why you cant write it as a decimal

OpenStudy (welshfella):

The line is an approximation anyway as the points are scattered

OpenStudy (readforlife):

Ohk.... So.... In my book they choose the number that passes through the y-intercept.... I'm kinda lost how you got the -5 when it seems like -1 passes through the y-intercept

OpenStudy (welshfella):

What we could tell the students is they can expect, on average, to gain about 1 point for every session they attend.

OpenStudy (welshfella):

Its -0,5 - about half way between the 0 and -1 on y axis

OpenStudy (readforlife):

Ok.

OpenStudy (readforlife):

So the equation for question 3 like you said is y = (7/6)x - 0.5. Ok

OpenStudy (welshfella):

I dont thikn theres only 1 answer to this y = (7/6) - 1 and y = (7/6)x are also acceptable ,

OpenStudy (welshfella):

Yes i think y = (7/6)x - 0.5 is a valid answer

OpenStudy (readforlife):

Ok. Thank you! So now for 4. do we solve the equation for part one of the question?

OpenStudy (welshfella):

well we've about covered number 4

OpenStudy (welshfella):

we can explain to attendants how to use the equation for predicting how many points they will probably get after attend ing 20 sessions:- number of points y = (7/6)*20 - 0.5 this comes to 22.8 points So we can round this up to 21 points.

OpenStudy (welshfella):

* i mean 23 points!

OpenStudy (welshfella):

For question 3 the x variables are the number of sessions and y variables are the number of points gained.

OpenStudy (readforlife):

Ok, thank you! And so sorry it took me so long. My wifi was acting up and I had to disconnect it and reload everything. Thank you so much @welshfella !!!